Sangria Topochico - The Capital Budgeting Decision In December 2012, María Guadalupe, the owner of Sidral Mundet Sol, had jusExhibit 1. Rates of Overweight and Obesity Perosns by Country (2010) Overweight Obesity Jupes (2005) Korea (2005) Swisd(2007)Sidral Mundet Sols products were sold only in small, independent grocery stores and convenience stores in Mexico. The firm aThe proposal Reading once again the executive summary of the report, María recalled what his father told him several times duoffering a longer collection period, letting the grocers pay in 45 days, instead of the normal 30 days. As far as accounts paAppendix 2. Sidral Mundet Sol—Balance Sheet as of December 31 (thousands of pesos) Assets 2009 2010 2011 Cash Account Receiva1. What are the background of the investment environment and the soda market in Mexico? 2. What are the relevant cash flows?
Answer #1

1. The investment environment was such that the company had not done introduced a product line in the last 5 years and launching a new line in the zero-calory seemed like a good idea. Because if we go by the fact the obese had tripled in Mexico and had the highest overweight rate in the world.

Talking about the Soda, it is known that the Mexicans loved their 'Soda Pop' and would drink it regularly due to lack of hygienic water. And the high per capita consumption (163l) resulted in obesity problems. In this category Coca-Cola, Pepsi-Cola, Dr. Pepper Snapple and Grupo Penafiel were the market leaders with 90% share and Coca-Cola being more dominant than the others. The revenues at the rate of 6.3% and consumption volumes grew at 4.5% between 2007-11. Though there was a dominant share of these players they commanded high prices and there was a gap in which the poorer people could be targeted with lower-priced soda.

2. The relevant cash flows in evaluating the project are : (only incremental cash flows) in pesos

  • Sales - 7,200,000 litres per year (600,000 * 12) @ 5 = 36,000,000 per year
  • Market Study - (5,000,000)
  • Capex - (50,000,000)
  • Depreciation - (9,200,000) per year
  • Resale Value - 4,000,000 at the end of 5th year

Raw material - (7,200,000) * 1.8 = 12,960,000 per year

labour cost - (180,000 * 12) = 2,160,000 per year

Energy cost = (50,000 * 12) = 600,000 per year

Administrative and General expenses = 300,000 per year

Accounting = (1% * 3,600,000) = 360,000 per year

Erosion = 800,000 per year

3. Erosion of existing product should be taken into consideration as the incremental cash flows to the firm decreases by that amount. Had that project not been undertaken by the company the firm would have earned the amount bu now it would have to forgo that amount and hence should be taken into consideration.

4. 2013 2014 2015 2016 2017 Revenue 36000000 36000000 36000000 36000000 36000000 less: production expenses Raw Material Labour E